# Measuring the Dunkelflaute: How (not) to analyze variable renewable energy shortage

### v1

@@ Abstract
As variable renewable energy sources increasingly gain importance in global energy systems, there is a growing interest in understanding periods of variable renewable energy shortage (“Dunkelflauten”). Defining, quantifying, and comparing such shortage events across different renewable generation technologies and locations presents a surprisingly intricate challenge. Various approaches exist in different bodies of literature, such as hydrology, wind and solar energy analysis, or energy system modeling. The subject of interest in previous analyses ranges from single technologies in specific locations to diverse technology portfolios across multiple regions, focusing either on supply from variable renewables or its mismatch with electricity demand. We provide an overview of methods for quantifying variable renewable energy shortage. We explain and critically discuss the merits and challenges of different approaches for defining and identifying shortage events and propose further methodological improvements for more accurate shortage determination. Additionally, we elaborate on comparability requirements for multi-technological and multi-regional energy shortage analysis. In doing so, we aim to contribute to unifying disparate methodologies, harmonizing terminologies, and providing guidance for future research.
@@
CSP
Centralized Solar Power
ENTSO-e
European Network of Transmission System Operators for Electricity
@@
fixed-duration MAZ
@@
fixed-duration MBT
LWP
low-wind-periods
@@
variable-duration MAZ
@@
variable-duration MBT
@@ 1 Introduction
Motivation: Dunkelflauten. Wind and solar power are cornerstone technologies for achieving climate-neutral energy systems around the world lee_ipcc_2023 . Because of their variable generation profiles, their system integration requires an increasing amount of temporal and spatial power sector flexibility. This includes, for example, different types of energy storage, demand-side flexibility, and geographical balancing denholm_grid_2011 ; rasmussen_storage_2012 ; schlachtberger_benefits_2017 ; roth_geographical_2023 ; he_sector_2021 ; kirchem_power_2023 . While various implications of renewable variability have been explored in the literature, attention has recently shifted to the occurrence of prolonged periods with extremely low wind and/or solar availability, also referred to as “Dunkelflauten”. This applies to both the research and the energy policy discourse in many countries dawkins_characterising_2020 ; huneke_f_kalte_2017 ; dhu_hintergrundpapier_2021 ; deutscher_bundestag_sicherstellung_2019 ; wood_go_2021 ; novacheck_evolving_2021 .
@@
Relevance of VRE shortage analysis. There is a growing interest in the questions as to how frequent and severe such shortage events may become in future energy systems heavily relying on variable renewable energy (VRE), and how to deal with them. Periods of VRE shortage will particularly require the use of long-duration energy storage dowling_role_2020 ; sepulveda_design_2021 as the potential for geographical balancing may be limited. This is because of the synoptic scale of these weather events, arising across thousands of kilometers brown_meteorology_2021 . However, a unique and generally accepted definition of VRE shortage events has not yet been established. This is because defining and quantifying VRE shortage events is a surprisingly intricate task that can be addressed from different angles, requiring a range of methodological choices and parameter assumptions. With this paper, we aim to provide comprehensive guidance for the analysis of VRE shortage conditions.
@@
Diverse approaches in VRE shortage analysis. VRE shortage events can be analyzed using different types of input data. One approach focusing on energy supply is to detect periods with low VRE availability, which can be either derived from measurements, modeled weather variables, or further processed recognizing technological assumptions and spatial distribution of the installed capacity. Many different terms have been used in the literature for such shortage events (Figure 1), e.g., “wind drought” potisomporn_extreme_2024 , “wind and solar resource drought” rinaldi_wind_2021 , or“energy production droughts” raynaud_energy_2018 in the literature. Alternatively, analyses may consider not only VRE generation potentials but also how these relate to electric load profiles. This can be done by analyzing residual load, sometimes also referred to as net load, which represents the net balance of electricity demand and VRE supply. Again, there is no consensus in the literature on how to term such shortage events (Figure 1), which, for example, have been referred to as “energy supply droughts” raynaud_energy_2018 , “energy deficits” ruhnau_storage_2022 , or “energy shortfall events” van_der_wiel_influence_2019 . For clarity, we propose the following terminology: “VRE droughts” for periods with low resource availability of a single or multiple VRE technologies, “ positive residual load (PRL) events” for periods where VRE supply falls short of electric load, and “VRE shortages” as a generic term that captures both concepts (Figure 1).
@@
generic term including both approaches
pt
@@
extreme low wind generation events Cannon et al. cannon_using_2015
low wind generation events Potisomporn & Vogel potisomporn_spatial_2022
low wind speed events Potisomporn et al. potisomporn_evaluating_2023
wind droughts Potisomporn et al. potisomporn_extreme_2024
hydrologic droughts Yevyevich yevyevich_objective_1967
streamflow drought Fleig et al. fleig_global_2006
extreme low wind events Abdelaziz et al. abdelaziz_assessing_2024
extended periods of low wind events Leahy & McKeogh leahy_persistence_2013
low wind speed events Patlakas et al. patlakas_low_2017
extreme low-wind-power events Ohlendorf & Schill ohlendorf_frequency_2020
energy production droughts Raynaud et al. raynaud_energy_2018
wind & solar resource droughts Rinaldi et al. rinaldi_wind_2021
wind & solar energy droughts Gangopadhyay et al. gangopadhyay_role_2022
low generation events Kies et al. kies_critical_2021
renewable energy droughts Kapica et al. kapica_potential_2024
renewable energy droughts Hu et al. hu_implications_2023
dark lulls Breyer et al. breyer_reflecting_2022
low wind & solar production events Kaspar et al. kaspar_climatological_2019
Dunkelflauten Mockert et al. mockert_meteorological_2023
Dunkelflaute Mayer et al. mayer_probabilistic_2023
dark doldrums Ohba et al. ohba_climatology_2022
energy production droughts Otero et al. otero_copula-based_2022
energy production droughts Allen & Otero allen_standardised_2023
wind and solar (power) droughts Brown et al. brown_meteorology_2021
wind and solar energy drought Bracken et al. bracken_standardized_2024
energy production drought Bracken et al. bracken_standardized_2024
low wind & solar production events van der Wiel et al. van_der_wiel_influence_2019
pt
@@
energy supply droughts Raynaud et al. raynaud_energy_2018
gaps between VRE supply & demand Rinaldi et al. rinaldi_wind_2021
Dunkelflaute in high demand period Mayer et al. mayer_probabilistic_2023
energy compound events Otero et al. otero_characterizing_2022
high residual load droughts Otero et al. otero_copula-based_2022
energy supply droughts Allen & Otero allen_standardised_2023
energy deficits Ruhnau & Qvist ruhnau_storage_2022
wind-drought-peak-demand Dawkins & Rushby dawkins_characterising_2020
electricity supply deficit Francois et al. francois_statistical_2022
low resource days Plain et al. plain_accounting_2019
energy supply drought Bracken et al. bracken_standardized_2024
energy shortfall events van der Wiel et al. van_der_wiel_influence_2019
@@
Figure 1: Proposal for a distinctive terminology and overview of previously used terms for VRE shortage conditions.
@@
Diverse scope of VRE shortage analyses. VRE shortage analyses can be done with various methods, ranging from stylized, assumption-based projections of future energy mixes to more sophisticated, optimizing energy system models. Further, the scope of VRE shortage analyses is broad, ranging from the assessment of individual renewable technologies in specific locations to optimized portfolios of generation technologies and demand flexibility across large balancing areas, encompassing multiple countries. Consequently, the definitions and identification methods for VRE shortage events should facilitate meaningful comparisons across various renewable generation technologies and regions.
@@
Literature: wind analysis. Different bodies of literature contribute to the analysis of VRE shortage events, primarily related to meteorology or energy research. One literature strand focuses on wind energy. Notably, wind droughts in the United Kingdom are well-studied. Cannon et al. cannon_using_2015 explore extreme onshore wind power events in the UK based on reanalysis data covering 33 years. Thresholds correspond to percentiles of the cumulative frequency distribution of investigated data. Potisomporn & Vogel potisomporn_spatial_2022 study the spatial and temporal correlation of low-wind-power events for offshore wind locations in the United Kingdom given a historical and projected distribution of offshore wind farms. Potisomporn et al. potisomporn_evaluating_2023 contrast low-wind speed events derived from ERA5 reanalysis data to measured data from 205 onshore and offshore wind locations in the United Kingdom. The same authors potisomporn_extreme_2024 compare different identification methods for low-wind events in the United Kingdom based on 72 years of ERA5 reanalysis data, deploying concepts for analyzing hydrological droughts yevyevich_objective_1967 ; fleig_global_2006 . Abdelaziz et al. abdelaziz_assessing_2024 assess offshore wind speed droughts in the United Kingdom. They identify consecutive days with daily average wind speed below a given threshold. There are related studies analyzing other regions. Leahy & McKeogh leahy_persistence_2013 and Patlakas et al. patlakas_low_2017 study the persistence of low-wind speeds for different locations in Ireland and the North Sea based on historical records. They search for events with wind speeds constantly below a threshold and for annual minimal values of moving average wind speeds over pre-defined intervals, applying extreme value theory. Thresholds are based on a range of exogenously set wind speeds. Ohlendorf & Schill ohlendorf_frequency_2020 investigate the duration and frequency of low onshore wind power events in Germany based on reanalysis data and exogenously set thresholds.
@@
Literature: VRE portfolio analysis. While the first VRE drought analyses focused on wind power, research interest in energy droughts of VRE portfolios that also include solar photovoltaics (PV) is growing. Raynaud et al. raynaud_energy_2018 investigate droughts for wind, solar, and hydro, both separately and combined as a portfolio, for 12 European regions and 30 years of data. They employ relative regional thresholds as fractions of the mean daily availability factor. Rinaldi et al. rinaldi_wind_2021 draw on Raynaud et al. raynaud_energy_2018 for identifying wind, solar, and wind-solar droughts in the Western U.S. for 39 years of data for different thresholds relative to mean daily generation. Gangopadhyay et al. gangopadhyay_role_2022 apply a similar approach to India, drawing on 5000 years of synthetic data. Kies et al. kies_critical_2021 compare seven VRE availability datasets for Europe. Using a composite VRE portfolio time series that combines wind and PV availability, they search for periods constantly below a range of exogenous thresholds. Kapica et al. kapica_potential_2024 compare energy droughts for solar PV and wind in Europe, either separately or as a combined portfolio, for historical and future weather conditions retrieved from eight regional climate models. Their identification method again draws on raynaud_energy_2018 with thresholds relative to seasonal and regional averages. Hu et al. hu_implications_2023 explore the risk of simultaneously occurring VRE droughts for European countries for single VRE technologies and a VRE portfolio, and compare it to droughts arising across all of Europe assuming perfect electricity exchange. Breyer et al. breyer_reflecting_2022 model decarbonization pathways for Europe until 2040 or 2050, using the weather year 2005. They screen their model outcomes for drought periods constantly below a range of thresholds relative to the maximum availability of wind and PV separately and combined for Germany and Europe. Kaspar et al. kaspar_climatological_2019 analyze wind and PV separately and combined for Germany based on reanalysis data. For given averaging intervals, they scan for moving averages below exogenously set thresholds. Mockert et al. mockert_meteorological_2023 also study VRE droughts in Germany in the context of prevailing weather regimes and lean on Kaspar et al. kaspar_climatological_2019 for drought identification. Mayer et al. mayer_probabilistic_2023 analyze concurrent PV and wind droughts in Hungary by identifying consecutive time steps with availability from both technologies below exogenous thresholds. Ohba et al. ohba_climatology_2022 link wind and solar droughts, both separately and combined as a portfolio, to four typical weather patterns arising in eastern Japan using an exogenously set threshold and neural network learning.
@@
Literature: residual load analysis. Finally, VRE shortage events are analyzed in the context of residual load. In addition to their VRE drought analysis, Raynaud et al. raynaud_energy_2018 investigate the imbalance of demand and VRE supply in Europe. Van der Wiel et al. van_der_wiel_influence_2019 test whether low wind and solar generation as well as high residual events are linked to large-scale weather regimes arising over Europe. Otero et al. otero_characterizing_2022 investigate periods with simultaneously low wind and solar generation and high demand across 27 European countries based on absolute thresholds. Otera et al. otero_copula-based_2022 use copula analysis to analyze VRE droughts based on a threshold relative to a ten percent fraction of its mean availability, and residual load events above the 90% percentile of its distribution in Europe based on ERA5 reanalysis. Allen & Otero allen_standardised_2023 introduce standardized VRE droughts indices derived from meteorological indicators that incorporate both VRE generation and electricity demand. They showcase the proposed method for droughts in Europe. Ruhnau & Qvist ruhnau_storage_2022 retrieve residual load time series from a model analysis of a least-cost decarbonized Germany power sector scenario and search for persistent positive residual load events. Dawkins & Rushby dawkins_characterising_2020 characterize wind droughts and residual load events that qualify as droughts in the United Kingdom based on percentile thresholds. Francois et al. francois_statistical_2022 explore positive residual load events (coined electricity deficits) for run-of-river hydropower and solar PV net of demand for Northern Italy. They deploy Complete-Time-Analysis, which is focused on return periods lasting from single hours up to several years. Based on the assumption of sufficient storage availability, Plain et al. plain_accounting_2019 identify low-solar periods using 21 years of satellite data to size microgrids power systems in Africa based on the mismatch between solar production and demand. Bracken et al. bracken_standardized_2024 characterize VRE droughts across the U.S. concerning frequency, duration, magnitude, and seasonality given the existing energy infrastructure in 2020. They investigate single VRE technologies, VRE portfolios, and positive residual load events.
@@
Contribution - harmonization. The review shows that different bodies of literature, e.g., hydrology, wind and solar energy, or energy system modeling, use a wide range of methodological approaches for defining and quantifying VRE shortage. Respective studies differ in terms of the technological, geographical, and temporal scope of the investigated system, based on partly inconsistent terminology. This could be a result of the disconnect of involved research fields. Communities at the intersections of these previously disconnected fields have established, bridging the gap between meteorology and energy research bloomfield_importance_2021 ; craig_overcoming_2022 . Yet, the literature lacks a structured overview and explanation of different methodological approaches for VRE shortage analyses using harmonized terminology. We aim to contribute to the literature by closing this gap.
@@
Contribution - VRE supply & residual load. We provide a methodological overview of different approaches for identifying VRE droughts across multiple regions and VRE technologies from the perspective of the energy system. We discuss possible definitions of VRE droughts and identification methods, focusing on renewable availability time series. As VRE potentials vary across technologies, space, and time, we elaborate on how to enable comparability across these dimensions. We further discuss how VRE drought analysis can be extended to also include electric load, i.e., how to evaluate extreme PRL. This allows for identifying compound low-VRE availability and high-demand events indicating the need for dispatchable capacity. Drawing on the proposed method for VRE availability time series, we discuss how to analyze PRL events and discuss essential points of criticism.

### v2

@@ Abstract
As variable renewable energy sources increasingly gain importance in global energy systems, there is a growing interest in understanding periods of variable renewable energy shortage (“Dunkelflauten”). Defining, quantifying, and comparing such shortage events across different renewable generation technologies and locations presents a surprisingly intricate challenge. Various methodological approaches exist in different bodies of literature, which have been applied to single technologies in specific locations or technology portfolios across multiple regions. We provide an overview of various methods for quantifying variable renewable energy shortage, focusing either on supply from variable renewables or its mismatch with electricity demand. We explain and critically discuss the merits and challenges of different approaches for defining and identifying shortage events and propose further methodological improvements for more accurate shortage determination. Additionally, we elaborate on comparability requirements for multi-technological and multi-regional energy shortage analysis. In doing so, we aim to contribute to unifying disparate methodologies, harmonizing terminologies, and providing guidance for future research.
@@
fixed-duration Mean-Above-Zero
@@
fixed-duration Mean-Below-Threshold
@@
variable-duration Mean-Above-Zero
@@
variable-duration Mean-Below-Threshold
@@ 1 Introduction
Wind and solar power are cornerstone technologies for achieving climate-neutral energy systems around the world [1]. Because of their variable generation profiles, their system integration requires an increasing amount of temporal and spatial power sector flexibility. This includes, for example, different types of energy storage, demand-side flexibility, and geographical balancing [2, 3, 4, 5, 6, 7]. While various implications of renewable variability have been explored in the literature, attention has recently shifted to the occurrence of prolonged periods with extremely low wind and/or solar availability, also referred to as “Dunkelflauten”. This applies to both the research and the energy policy discourse in many countries [8, 9, 10, 11, 12, 13].
@@
There is a growing interest in the questions as to how frequent and severe such shortage events may become in future energy systems heavily relying on variable renewable energy (VRE), and how to deal with them. Periods of VRE shortage will particularly require the use of long-duration energy storage [14, 15] as the potential for geographical balancing may be limited. This is because of the synoptic scale of these weather events, arising across thousands of kilometers [16]. However, a unique and generally accepted definition of VRE shortage events has not yet been established. This is because defining and quantifying such events is a surprisingly intricate task that can be addressed from different angles, requiring a range of methodological choices and parameter assumptions. With this paper, we aim to provide comprehensive guidance for the analysis of variable renewable energy shortage.
VRE shortage can be analyzed using different types of input data. One approach that focuses on energy supply is to detect periods with low renewable availability. The latter can be derived from measurements, modeled weather variables, or renewable availability time series, recognizing technological assumptions and spatial distribution of the installed capacity. Shortage can be detected by identifying concrete periods with availability below a given threshold. For this, many different terms have been used in the literature (Figure 1), e.g., “wind drought” [17], “wind and solar resource drought” [18], or “energy production droughts” [19] in the literature. An alternative is to compute deviations of renewable availability from a reference profile over a user-specific time horizon. The latter has been termed “seasonal variability”, “weather variability”, or “wind droughts” [20], as well as “cumulative anomaly of a renewable resource” [21].
@@
Another approach is to consider not only VRE generation potentials, but also how these relate to electric load profiles. This can be done by analyzing residual load, sometimes also referred to as net load, which represents the net balance of electricity demand and VRE supply. Again, there is no consensus in the literature on how to term such shortage events (Figure 1), which, for example, have been referred to as “energy supply droughts” [19], “energy deficits” [22], or “energy shortfall events” [23]. For clarity, we propose the following terminology: “variable renewable energy droughts” for periods with low resource availability of a single or multiple VRE technologies, “variable renewable energy anomaly” for cumulative deviations of VRE resource availability from some mean reference, and “ positive residual load (PRL) events” for periods where variable renewable supply falls short of electric load, and “variable renewable energy shortage” as a generic term that captures these concepts (Figure 1). Note that VRE drought and PRL events refer to concrete periods and do not necessarily occur at the same time. While both concepts characterize variable renewable energy shortage, PRL events are additionally affected by the timing of coinciding high-demand periods. In contrast, VRE anomalies are measured over user-specific time horizons.
@@
generic term
@@
extreme low wind generation events Cannon et al. [24]
low wind generation events Potisomporn & Vogel [25]
low wind speed events Potisomporn et al. [26]
wind droughts Potisomporn et al. [17]
hydrologic droughts Yevyevich [27]
streamflow droughts Fleig et al. [28]
extreme low wind events Abdelaziz et al. [29]
extended periods of low wind events Leahy & McKeogh [30]
low wind speed events Patlakas et al. [31]
extreme low-wind-power events Ohlendorf & Schill [32]
energy production droughts Raynaud et al. [19]
wind & solar resource droughts Rinaldi et al. [18]
wind & solar energy droughts Gangopadhyay et al. [33]
low generation events Kies et al. [34]
renewable energy droughts Kapica et al. [35]
renewable energy droughts Hu et al. [36]
dark lulls Breyer et al. [37]
low wind & solar production events Kaspar et al. [38]
Dunkelflauten Mockert et al. [39]
Dunkelflauten Mayer et al. [40]
dark doldrums Ohba et al. [41]
energy production droughts Otero et al. [42]
energy production droughts Allen & Otero [43]
wind and solar (power) droughts Brown et al. [16]
wind and solar energy droughts Bracken et al. [44]
energy production droughts Bracken et al. [44]
low wind & solar production events van der Wiel et al. [23]
variable renewable energy anomaly
focusing on resource availability,
identification of deviations from reference availability,
considering VRE availability time series only
seasonal variability, weather variability, Antonini et al. [20]
and wind droughts
climatological renewable energy Stoop et al. [21]
deviation index (CREDI)
@@
energy supply droughts Raynaud et al. [19]
gaps between VRE supply & demand Rinaldi et al. [18]
Dunkelflaute in high demand period Mayer et al. [40]
energy compound events Otero et al. [45]
high residual load droughts Otero et al. [42]
energy supply droughts Allen & Otero [43]
energy deficits Ruhnau & Qvist [22]
wind-drought-peak-demand Dawkins & Rushby [8]
electricity supply deficits Francois et al. [46]
low resource days Plain et al. [47]
energy supply droughts Bracken et al. [44]
energy shortfall events van der Wiel et al. [23]
@@
Figure 1: Proposal for a distinctive terminology and overview of previously used terms in two major and one emerging strand of VRE shortage analysis.
@@
VRE shortage analyses can be carried out with various methods, ranging from stylized, assumption-based projections of future energy mixes to more sophisticated, optimizing energy system models. Further, the scope of variable renewable energy shortage analyses is broad, ranging from the assessment of individual renewable technologies in specific locations to optimized portfolios of generation technologies and demand flexibility across large balancing areas, encompassing multiple countries. Consequently, the definitions and identification methods for VRE shortage events should facilitate meaningful comparisons across various renewable generation technologies and regions.
@@
Different bodies of literature contribute to the analysis of VRE shortage events, primarily related to meteorology or energy research. One literature strand focuses on wind energy. Notably, wind droughts in the United Kingdom are well-studied. Cannon et al. [24] explore extreme onshore wind power events in the UK based on reanalysis data covering 33 years. Thresholds correspond to percentiles of the cumulative frequency distribution of investigated data. Potisomporn & Vogel [25] study the spatial and temporal correlation of low-wind-power events for offshore wind locations in the United Kingdom given a historical and projected distribution of offshore wind farms. Potisomporn et al. [26] contrast low-wind speed events derived from ERA5 reanalysis data to measured data from 205 onshore and offshore wind locations in the United Kingdom. The same authors [17] compare different identification methods for low-wind events in the United Kingdom based on 72 years of ERA5 reanalysis data, deploying concepts for analyzing hydrological droughts [27, 28]. Abdelaziz et al. [29] assess offshore wind speed droughts in the United Kingdom. They identify consecutive days with daily average wind speed below a given threshold. There are related studies analyzing other regions. Leahy & McKeogh [30] and Patlakas et al. [31] study the persistence of low-wind speeds for different locations in Ireland and the North Sea based on historical records. They search for events with wind speeds constantly below a threshold and for annual minimal values of moving average wind speeds over pre-defined intervals, applying extreme value theory. Thresholds are based on a range of exogenously set wind speeds. Ohlendorf & Schill [32] investigate the duration and frequency of low onshore wind power events in Germany based on reanalysis data and exogenously set thresholds.
@@
While the first VRE drought analyses focused on wind power, research interest in droughts of variable renewable energy portfolios that also include solar photovoltaics (PV) is growing. Raynaud et al. [19] investigate droughts for wind, solar, and hydro, both separately and combined as a portfolio, for 12 European regions and 30 years of data. They employ relative regional thresholds as fractions of the mean daily availability factor. Rinaldi et al. [18] draw on Raynaud et al. [19] for identifying wind, solar, and wind-solar droughts in the Western U.S. for 39 years of data for different thresholds relative to mean daily generation. Gangopadhyay et al. [33] apply a similar approach to India, drawing on 5000 years of synthetic data. Kies et al. [34] compare seven VRE availability datasets for Europe. Using a composite VRE portfolio time series that combines wind and PV availability, they search for periods constantly below a range of exogenous thresholds. Kapica et al. [35] compare energy droughts for solar PV and wind in Europe, either separately or as a combined portfolio, for historical and future weather conditions retrieved from eight regional climate models. Their identification method again draws on [19] with thresholds relative to seasonal and regional averages. Hu et al. [36] explore the risk of simultaneously occurring VRE droughts for European countries for both single technologies and a VRE portfolio, and compare it to droughts arising across all of Europe assuming perfect electricity exchange. Breyer et al. [37] model decarbonization pathways for Europe until 2040 or 2050, using the weather year 2005. They screen their model outcomes for drought periods constantly below a range of thresholds relative to the maximum availability of wind and PV separately and combined for Germany and Europe. Kaspar et al. [38] analyze wind and PV separately and combined for Germany based on reanalysis data. For given averaging intervals, they scan for moving averages below exogenously set thresholds. Mockert et al. [39] also study VRE droughts in Germany in the context of prevailing weather regimes and lean on Kaspar et al. [38] for drought identification. Mayer et al. [40] analyze concurrent PV and wind droughts in Hungary by identifying consecutive time steps with availability from both technologies below exogenous thresholds. Ohba et al. [41] link wind and solar droughts, both separately and combined as a portfolio, to four typical weather patterns arising in eastern Japan using an exogenously set threshold and neural network learning.
@@
In another strand of the literature, VRE shortage events are analyzed in the context of residual load. In addition to their VRE drought analysis, Raynaud et al. [19] investigate the imbalance of demand and VRE supply in Europe. Van der Wiel et al. [23] test whether low wind and solar generation as well as high residual events are linked to large-scale weather regimes arising over Europe. Otero et al. [45] investigate periods with simultaneously low wind and solar generation and high demand across 27 European countries based on absolute thresholds. Otera et al. [42] use copula analysis to analyze VRE droughts based on a threshold relative to a ten percent fraction of its mean availability, and residual load events above the 90% percentile of its distribution in Europe based on ERA5 reanalysis. Allen & Otero [43] introduce standardized VRE droughts indices derived from meteorological indicators that incorporate both VRE generation and electricity demand. They showcase the proposed method for droughts in Europe. Ruhnau & Qvist [22] retrieve residual load time series from a model analysis of a least-cost decarbonized Germany power sector scenario and search for persistent positive residual load events. Dawkins & Rushby [8] characterize wind droughts and residual load events that qualify as droughts in the United Kingdom based on percentile thresholds. Francois et al. [46] explore positive residual load events (coined electricity deficits) for run-of-river hydropower and solar PV net of demand for Northern Italy. They deploy Complete-Time-Analysis, which is focused on return periods lasting from single hours up to several years. Based on the assumption of sufficient storage availability, Plain et al. [47] identify low-solar periods using 21 years of satellite data to size microgrids power systems in Africa based on the mismatch between solar production and demand. Bracken et al. [44] characterize VRE droughts across the U.S. concerning frequency, duration, magnitude, and seasonality given the existing energy infrastructure in 2020. They investigate single VRE technologies, VRE portfolios, and positive residual load events.
Recently, another literature strand has been emerging in which cumulative VRE anomalies from some reference over a given time horizon are identified. Antonini et al. [20] identify wind energy droughts on a global scale as deviation of hourly wind power density from its multidecadal, climatological mean hourly profile. They identify world well-suited wind regions characterized by high wind power density, low inter- and intra-annual variability, and less pronounced droughts. Stoop et al. [21] follow a similar approach and introduce a climatological renewable energy deviation index (CREDI), which quantifies cumulative anomalies of renewable resources over different timescales, ranging from hours to decades. While this method is not limited to analyses of renewable energy shortage events, the authors present an exemplary application for 8-day Dunkelflaute events in the paper. The approaches in this emerging literature strand avoid some of the methodological challenges of VRE availability and PRL event analysis that will be discussed below. Yet, the choice of the time horizon remains user-specific. Further, it is not clear how the chosen reference profiles that are used to compute renewable anomalies relate to energy system needs.
Alternatively, Grochowicz et al. [48] detect power system stress events in Europe based on electricity shadow prices derived from a power sector model. They consider both large-scale weather patterns over 40 years of re-analysis data and power sector dynamics. For each year, they identify system-defining events that may last up to two weeks and account for a considerable share of total system costs. Subsequently, they classify the mid-troposphere dynamics into weather regimes and cluster near-surface meteorological conditions present during these events. System-defining events occur between November and February and last between two and 13 days. The analysis also takes into account the shadow prices of transmission capacity constraints and hourly storage energy constraints. Importantly, their findings are only applicable to the particular model setting and rely on a range of model-inherent parameter assumptions.
@@
The review shows that different bodies of literature, e.g., hydrology, wind and solar energy, or energy system modeling, use a wide range of methodological approaches for defining and quantifying VRE shortage. While alternative approaches are emerging, the vast majority of the reviewed studies use threshold-based methods for analyzing variable renewable energy drought and positive residual load events (in the latter case, using a zero threshold). Additionally, existing studies differ in terms of the technological, geographical, and temporal scope of the investigated system, based on partly inconsistent terminology. This could be a result of the disconnect of involved research fields. Communities at the intersections of these previously disconnected fields have established, bridging the gap between meteorology and energy research [49, 50]. Yet, the literature lacks a structured overview and explanation of different methodological approaches for VRE shortage analyses using harmonized terminology. We aim to contribute to the literature by closing this gap.
@@
We provide a structured overview of different approaches for identifying variable renewable energy droughts across multiple regions and technologies from the perspective of the energy system. Based on the literature review, we focus our analysis on threshold-based methods, including variable renewable energy droughts and positive residual load events. We discuss possible definitions of VRE droughts and identification methods, focusing on renewable availability time series. As variable renewable energy potentials vary across technologies, space, and time, we elaborate on how to enable comparability across these dimensions, which has not yet been considered in the literature. We further discuss how VRE drought analysis can be extended to also include electric load, i.e., how to evaluate extreme PRL. This allows for identifying compound low-VRE availability and high-demand events indicating the need for dispatchable capacity. Drawing on the proposed method for VRE availability time series, we discuss how to analyze PRL events and discuss essential points of criticism.